OpenStudy (anonymous):

One number is 5 more than twice another and their sum is 23. Which of the following system of equations represents the word problem? x + y = 23 and y = 2x + 5 y = x + 5 and y + 2(x + 5) = 23 y = 2(x + 5) and x + y = 23

OpenStudy (anonymous):

the first one you wrote, or x + y = 23 and y = 2x + 5 represents this word problem. Its easier if you break it down. One number (y) is (=) 5 more than twice another number (2x + 5). So then you get y = 2x + 5. Then, these 2 numbers have a sum of 23, or in other words, are added together to get 23. So, the two numbers (x and y) are added together to get (=) 23. so from there you get x + y= 23. this is why the first choice you wrote is correct. hope i helped!